OpenStudy (anonymous):

what is the simplified form of 20x^6/15 y^5 times 5y^2/6x^4

OpenStudy (anonymous):

\[\frac{ 20x ^{6} }{ 15y ^{2}} \times \frac{ 5y ^{2} }{ 6x ^{4} }\]

OpenStudy (luigi0210):